OpenStudy (anonymous):

how can i simplify (14+7√(2))/7 to 2+√(2)?

OpenStudy (anonymous):

is that two in double parenthesis?

OpenStudy (anonymous):

is it the denom of 7 divided by the num of 14 and 7?

OpenStudy (anonymous):

either or it doesnt matter

OpenStudy (anonymous):

1st because its a square root and 2nd because its closing the equation

OpenStudy (anonymous):

\[14+7\sqrt{2}/7+2\sqrt{2}\]

OpenStudy (anonymous):

\[14+7\sqrt{2}/7\]

OpenStudy (anonymous):

simplify that to \[2\sqrt{2}\]

OpenStudy (anonymous):

21 ( sq root) 2/ 9 (sq. root) 2 - thats how i would write it out

OpenStudy (anonymous):

the 7 goes to the 2

OpenStudy (anonymous):

the answer is \[2\sqrt{2}\], but i didnt know how to get it

OpenStudy (anonymous):

i think i figured it out... by dividing the denom(7) by the num(14,7)

OpenStudy (anonymous):

was that one two in double parenthesis then it gets multiplied

OpenStudy (ajprincess):

Is this ur question? \[\frac{ 14+7\sqrt2 }{ 7 }\]

OpenStudy (anonymous):